OpenStudy (anonymous):

the length of a rectangle is 7 meters more than the width. the area is 60 square meters. find the length and width.

jimthompson5910 (jim_thompson5910):

let x be the width

jimthompson5910 (jim_thompson5910):

it says that "the length of a rectangle is 7 meters more than the width", so the length is x+7

jimthompson5910 (jim_thompson5910):

Area = Length * Width 60 = (x+7)*x 60 = x^2 + 7x 0 = x^2 + 7x - 60 x^2 + 7x - 60 = 0 I'll let you finish

OpenStudy (anonymous):

width = 5 and length is 12

jimthompson5910 (jim_thompson5910):

PI3.14 please don't give out the answers only

OpenStudy (anonymous):

let the width be x so that the length will be 7 + x. now the area is x * (7 + x) = 60 that is x^2 + 7x - 60 = 0

OpenStudy (anonymous):

now solve for x and the positive value will be the width